If the terminal speed of a sphere of gold (density = 19.5 kg/m3) is 0.2 m/s in a viscous liquid (density = 1.5 kg/m3), find the terminal speed of a sphere of silver (density = 10.5 kg/m3) of the  same size in the same liquid
(a) 0.4m/s
(b) 0.133m/s
(c) 0.1 m/s
(d) 0.2 m/s

Grade:upto college level

2 Answers

Navjyot Kalra
9 years ago
(c) Terminal velocity, vT = 2r2(d1 – d2 )g
vT2/0.2 = (10.5 – 1.5)/(19.5 – 1.5) \Rightarrowv2 = 0.2 * 9/18 = 0.1 m/s
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Terminal velocity, vT = 2r2(d1 – d2 )g
vT2/0.2 = (10.5 – 1.5)/(19.5 – 1.5)
vT­2 = 0.2 * 9/18 = 0.1 m/s

Thanks and Regards